# SpeechAct: Towards Generating Whole-body Motion from Speech

### v3

@@ 4.4 User Study
We evaluate the percentage of each method considered to be ranked first in different perspectives. Fig. 8 shows the statistical results. it is evident that more than 70% of users believe that our method generates more realistic results, which indicates that our hybrid point representation can produce more smooth and reasonable results. Besides, compared to other methods, more than 70% of users think that our method produces results more in line with the audio rhythm, and 68% of users believe that our method can generate more diverse motion. These results demonstrate that our method surpasses other methods in terms of diversity and alignment with audio signals, which suggests the effectiveness of our hybrid point representation and our contrastive motion learning method.
@@
Figure 10: Visualization of generalization conditioned on music inputs.
@@ 4.5 Generalization
Our model generates reasonable and diverse motions from only audio signals. Therefore, even with training with English data of BEAT2 dataset [42], our model can be generalized to other languages. Fig. 9 shows the qualitative results. We can see that with the specific rhythm, our model can generate corresponding motions.
Moreover, our model can also be generalized to music inputs. The qualitative results can be found in Fig. 10. We can see that when the music is soothing, it generates smooth changes in motion (first sample), and when the music becomes intense, our method can generate movements corresponding to the rhythm (second sample). This means our model can generate different motions according to the change in the music. Besides, this also indicates our hybrid point representation and contrastive motion learning method can be used in other motion generation tasks, e.g., text-driven motion generation and music-driven motion generation. The dynamic results can be found in the supplementary video.
Figure 11: Avatar animation results. Given a speech, the generated motion can be used to animate reconstructed avatars.
@@ 4.6 Application
Thanks to our hybrid point representation, our model can output SMPLX parameters to support many applications. In this section, we obtain an avatar reconstructed from [58], and animate it using the outputs of our model. By doing so, we can get animated avatars directly using speech, which can be used in VR/AR, virtual live, and human-computer interaction. Fig. 11 presents the avatar animation results. It can be seen that the generated motion can be applied to avatar animation, and can obtain promising results. The dynamic results can be found in the supplementary video.
@@ 5 Conclusions and Discussions
Conclusion. In this paper, we propose a novel framework, named SpeechAct, with a face generator and a body generator for whole-body motion generation from speech. The face generator is used to generate deterministic results using an encode-decoder architecture. For the body generator, we propose a hybrid point representation for body and hand to constrain the global surface and capture local details for 3D body, which can achieve accurate yet continuous results. Based on our representation, we design a two-stage model with a novel contrastive motion learning method to achieve diverse body and hand motions. Experimental results demonstrate that our model can generate reasonable and diverse human motion. We give some promising results for generalization across other languages and music inputs. Moreover, an application of speech-driven avatar animation is given to show the potential of VR/AR and human-computer interaction.
Limitation and Future Work. To achieve natural and reasonable results, we estimate the global position using a residual estimation manner. Specifically, we estimate position residuals and add previous residuals to obtain the current position, which can help to generate smooth results. However, for long motion generation, the position error will be added and lead to a sliding effect for some specific cases. The visualization of some samples can be found in the supplemental document. In the future, we will consider more temporal constraints to generate more smooth results. Besides, our model can generate diverse results with distinctive styles. However, we only use the speaker embedding as the style condition, which limits the effectiveness of controllability. In the future, we will add more control inputs, e.g., expression and text, as conditional inputs to achieve more controllable human motion synthesis.
@@ Acknowledgements
This work was supported in part by National Natural Science Foundation of China (62122058 and 62171317), National Key R&\&&D Program of China (2023YFC3082100), and Science Fund for Distinguished Young Scholars of Tianjin (No. 22JCJQJC00040). We are grateful to Associate Editor and anonymous reviewers for their help in improving this paper.

### v5

@@ 4.4 User Study
We evaluate the percentage of each method considered to be ranked first in different perspectives. Fig. 10 shows the statistical results. it is evident that more than 70% of users believe that our method generates more realistic results, which indicates that our hybrid point representation can produce more smooth and natural results. Besides, compared to other methods, more than 70% of users think that our method produces results more in line with the audio rhythm, and 68% of users believe that our method can generate more diverse motion. We also conducted a significance test between our method and the other two methods, which indicates that our method shows statistically significant differences compared to the other two methods. These results demonstrate that our method surpasses other methods in terms of diversity and alignment with audio signals, which suggests the effectiveness of our hybrid point representation and our contrastive motion learning method. Details of the user study and the significance test can be found in the supplemental document.
@@ 4.5 Generalization
Our model generates natural and diverse motions from only audio signals. Therefore, even with training with English data of BEAT2 dataset [45], our model can be generalized to other languages. Fig. 11 shows the qualitative results. We can see that with the specific rhythm, our model can generate corresponding motions.
Figure 12: Avatar animation results. Given a speech audio, the generated motion can be used to animate reconstructed avatars.
@@ 4.6 Application
Thanks to our hybrid point representation, our model can output SMPLX parameters to support many applications. In this section, we obtain an avatar reconstructed from [68], and animate it using the outputs of our model. By doing so, we can get animated avatars directly using speech, which can be used in VR/AR, virtual live, and human-computer interaction. Since the expression basis of the facial model, i.e., ARKit, used in AvatarRex differs from ours, i.e., SMPLX, we retrained our model using the expression format of ARKit [69]. As a result, with the output motions generated by our method, we can directly animate the avatar reconstructed using AvatarRex. To showcase additional results, we adopt the recent avatar reconstruction method ExAvatar [70] to produce animation results based on our generated motion sequence. Notably, we can directly use the first 50 dimensions of our expression parameters as the expression parameters for ExAvatar, which ensures compatibility with our method without the need for retraining a new model. Fig. 12 presents the avatar animation results reconstructed from AvatarRex (the first row) and ExAvatar (the second row). It can be seen that the generated motion can be applied to avatar animation, and can obtain promising results. The dynamic results can be found in the supplementary video.
@@ 5 Conclusions and Discussions
Conclusion. In this paper, we propose a novel framework, named SpeechAct, with a face generator and a body generator for whole-body motion generation from speech audio. The face generator is used to generate deterministic results using an encode-decoder architecture. For the body generator, we propose a hybrid point representation for body and hand to constrain the global surface and capture local details for 3D body, which can achieve accurate yet continuous results. Based on our representation, we design a two-stage model with a novel contrastive motion learning method to achieve diverse body and hand motions. Experimental results demonstrate that our model can generate natural and diverse human motion. We give some promising results for generalization across other languages. Moreover, an application of speech-driven avatar animation is given to show the potential of VR/AR and human-computer interaction.
Limitation and Future Work. To achieve natural results, we estimate the global position using a residual estimation manner. Specifically, we estimate position residuals and add previous residuals to obtain the current position, which can help to generate smooth results. However, for long motion generation, the position error will be added and lead to a sliding effect for some specific cases. The visualization and discussion of some samples can be found in the supplemental document. In the future, we will consider more temporal constraints to generate more smooth results. Besides, our model can generate diverse results with distinctive styles. However, we only use the speaker embedding as the style condition, which limits the effectiveness of controllability. In the future, we will add more control inputs, e.g., text, as conditional inputs, and leverage the speaker information to achieve more controllable human motion synthesis.
@@ Acknowledgements
This work was supported in part by National Key R&\&&D Program of China (2023YFC3082100), National Natural Science Foundation of China (62122058 and 62171317), and Science Fund for Distinguished Young Scholars of Tianjin (No. 22JCJQJC00040). We are grateful to Associate Editor and anonymous reviewers for their help in improving this paper.